40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 183     Entry time: Mon May 12 22:43:02 2014
Author: Koji 
Type: Optics 
Category: Characterization 
Subject: Measured FSR/TSM of the OMC cavity 

Data analysis of the FSR/TSM measruement last week.

1. FSR was measured with "the golden arches" technique.

FSR = 263.0686 MHz +/- 900Hz

Lcav = 1.1396 m --> 7.6 mm too long! (nominal 1.132m)

2. Transverse mode spacings for the vertical and horizontal modes were measured.

TMS/FSR = 0.219366 (V) / 0.220230 (H) (Predicted value with the current cavity length 0.2196/0.2202 very close!)

We want to make this to be ~0.219 (~3% less)

With the current parameters, the 19th-order lower sideband make the coincident resonance.

Attachment 1: Cav_scan_response_Pitch.pdf  205 kB  | Hide | Hide all
Cav_scan_response_Pitch.pdf
Attachment 2: Cav_scan_response_Yaw.pdf  204 kB  | Hide | Hide all
Cav_scan_response_Yaw.pdf
ELOG V3.1.3-